# EVM-Fusion: An Explainable Vision Mamba Architecture with Neural Algorithmic Fusion

### v3

@@ 5 Discussion
Through the above analysis, the clinical application of the model can be enhanced. By analyzing these heat maps, if the model’s paths do not pay enough attention to the potential lesion area, even doctors who know nothing about the model principle can know that the output is unreliable. On the contrary, if both models pay attention to the organs that need attention, especially the U-Net path, then even if the doctor does not believe the output of the model, he can focus on the places that U-Net pays attention to and overlaps with the organs, because those places are likely to be the location of the lesions.
@@ 6 Conclusion
In this paper, I have introduced EVM-Fusion, a novel deep learning architecture designed to address the intricate demands of multi-organ medical image classification by synergizing advanced feature extraction, dynamic fusion, and intrinsic explainability. My approach leverages a multipath framework where Vision Mamba modules enhance features from DenseNet and U-Net backbones, complemented by a dedicated path for traditional handcrafted features. The core innovation lies in our two-stage fusion strategy, particularly the Neural Algorithmic Fusion (NAF) block, which, preceded by cross-modal attention, learns an iterative, adaptive algorithm to optimally combine diverse feature streams. This NAF mechanism, incorporating a controller that observes previous fusion states and utilizes residual connections, represents a significant departure from static fusion techniques.
Experimental validation on a challenging 9-class dataset, encompassing X-rays and MRIs from multiple organ systems, demonstrated the efficacy of EVM-Fusion, achieving excellent classification accuracy of 99.75% on the test set. The detailed XAI analysis, presented in Section 5 (Discussion), further elucidated the model’s behavior, showcasing its ability to dynamically focus on salient image regions, adaptively weigh features from different paths based on input characteristics (as evidenced by varying cross-modal attention), and leverage specific textural and structural cues identified by both deep and traditional feature sets. This dynamic interplay, particularly the adaptive path weighting via cross-modal attention feeding into the NAF block, likely contributes to the model’s robustness and superior performance compared to architectures relying on simpler or static fusion.

### v4

@@ 5 Discussion
Through the above analysis, the clinical application of the model can be enhanced. By analyzing these heat maps, if the model’s paths do not pay enough attention to the potential lesion area, even doctors who know nothing about the model’s principle can know that the output is unreliable. On the contrary, if both models pay attention to the organs that need attention, especially the U-Net path, then even if the doctor does not believe the output of the model, he can focus on the places that U-Net pays attention to and overlaps with the organs, because those places are likely to be the locations of the lesions.
@@ 6 Conclusion
This paper introduced EVM-Fusion. It is a synergistic framework designed for the clinical task of rapid, holistic medical image triage. We started by identifying the specific challenges of this task. These challenges include the need to process both local and global information and the difficulty of fusing heterogeneous features. We then presented a solution based on a Mamba-enhanced multi-path design and a learnable algorithmic fusion strategy.
Our experiments provided strong validation for this approach. The extensive ablation studies showed that every component of our framework is essential. Their combination leads to performance that is greater than the sum of the parts. Our model also outperformed a wide range of powerful, single-paradigm SOTA models. This shows the practical benefits of our synergistic design. Finally, the built-in XAI tools offer a necessary layer of transparency, which is crucial for building trustworthy AI solutions for clinical use.